Select the option for "?" that continues the pattern in each question.
3, -6, 12, 4, 20, ?
13
14
6
16
it's for a caliper test at my office. I have no idea why nor do I have any idea how to solve this...

Answer by math_helper(2461)



3, -6, 12, 4, 20, ?
Pattern seems to be 3-9 = -6, -6+18 = 12, 12-8 = 4, 4+16 = 20, ...
So the additions are, starting with 3: -9, +18, -8, +16, next is -7 (?)
The continuation I see is 20-7 = 13